jabez177 (jabez177):

MEDAL IF U CAN HELP PLZZZZZZZZ What is 18 divided by 2/3 in the lowest terms. I think it is 27 PLZ CHECK MY WORK

jabez177 (jabez177):

@Samus_Aran

OpenStudy (sreekar369):

u r correct

jabez177 (jabez177):

OK THNX

OpenStudy (sreekar369):

18/2/3=18*3/2=27

OpenStudy (mrbaconqwerty):

u r right @jabez177

jabez177 (jabez177):

THAT IS WHAT I DID

jabez177 (jabez177):

SO IT GOT ME 27 AND THNX

OpenStudy (anonymous):

no it 27.027

jabez177 (jabez177):

no it is 27 there is no 27.027

OpenStudy (anonymous):

ya i did it on a calculator

jabez177 (jabez177):

ok

OpenStudy (mrbaconqwerty):

i did it on my calculator, and got 27

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

same

jabez177 (jabez177):

and u 2 can debate i gtg peace \ /

OpenStudy (anonymous):

ya but 2/3 is 0.666 so 18/0.666=27.027

jabez177 (jabez177):

0r 4

OpenStudy (anonymous):

later gator XD

jabez177 (jabez177):

so c ya dudes and madam XD

OpenStudy (anonymous):

hay u gona play tanki today

jabez177 (jabez177):

yeah later ontry and add this gut he is low rank and my bro and i use it cuz i am helping him rank up Theycallmejuan_TURBO just like that

OpenStudy (anonymous):

ok i will later cya i gtg i write it down

jabez177 (jabez177):

ok